Mary's Mugs produces and sells various types of ceramic mugs. The business began operations on January 1, year 1, and its costs incurred during the year include these:
  

Variable costs (based on mugs produced):
Direct materials cost $ 2,100
Direct manufacturing labor costs 22,770
Indirect manufacturing costs 1,070
Administration and marketing 2,020
Fixed costs:
Administration and marketing costs 12,000
Indirect manufacturing costs 4,160


On December 31, year 1, direct materials inventory consisted of 2,100 pounds of material. Production in that year was 14,000 mugs. All prices and unit variable costs remained constant during the year. Sales revenues for year 1 was $43,450. Finished goods inventory was $6,450 on December 31, year 1. Each finished mug requires 0.4 pounds of material. (Do not round intermediate calculations.)
  

Required:

a. Compute the direct materials inventory cost, December 31, year 1.

  

b. Compute the finished goods ending inventory in units on December 31, year 1. (Do not round intermediate calculations.)

  

c. Compute the selling price per unit. (Round your answer to 2 decimal places.)

d. Compute the operating profit (loss) for year 1.
Answer #1
a. Compute the direct materials inventory cost, December 31, year 1.
Direct Material cost per unit = Direct material cost / units produced = $2100/14000 mugs = $0.14 per mug
Direct material used per mug = 0.40 pounds
Direct material cost per pound = $0.14/ 0.40 = $0.35 per round
Direct material inventory = 2100 Pound * $0.35 = $735
b. Compute the finished goods ending inventory in units on December 31, year 1.
Finished Goods inventory (in units) = Finished goods inventory / manufacturing cost per unit
Manufacturing cost per unit = (Direct material + Direct Labour + Indirect manufacturing cost)/Units Produced
= ($2100+$22,770+$1,070+$4,160)/14000 = $2.15 per unit
Finished Goods inventory (in unit) :
Year 1 = $6,450/$2.15 = 3000 units
c. Compute the selling price per unit.
Selling price per unit = Revenues / units sold
Units sold = Units produced - units in the ending finished goods inventory = 14000-3000 = 11000
Selling price per unit = $43,450/11000 = $3.95
d. Compute the operating profit (loss) for year 1
Operating income for the year :
Revenues (a) $43,450
Cost of goods sold (11000*$2.15) (b) $23,650
Gross Margin(c = a-b) $19,800
Less marketing and admin cost
   Variable cost $2,020
Fixed cost $12,000 $14,020
Operating Profit $5,780